# Observing structural disorder induced interacting topological phase in an atom array

### v2

@@ Abstract
Figure 3: Observation of structural disorder induced many-body average SPT phase. (A) Schematics of the experimental sequence for many-body ground state preparation and detection. Initially, all atoms are excited into the Rydberg $p$ state (top left panel). Focused pinning laser beams (in a red bowtie shape) are applied to atoms in the upper sub-chain, and a microwave field is swept across resonance to produce a Néel state (right panel). Adiabatic evolution subsequently prepares the ground state of the Hamiltonian $\hat{H}$ in a fixed particle number subspace. (B) Energy spectra at different particle numbers for a regular lattice and (C) disordered lattice ($W=0.4{d}$), averaged over $15$ random realizations with a zoomed-in view near the ground states. The four-fold degeneracy of the ground state is slightly lifted due to the van der Waals interactions. The average inversion symmetry only protects the two-fold ground state degeneracy in the half-filling subspace (see the SI). (D) Measured site occupancy of the ground state for the regular lattice compared to (E) the disordered lattice near half-fillings of $N-1$ (left), $N$ (right), and $N+1$ (below) particles. The measured results are postselected for specific particle fillings. Diagonal crosses represent numerical results incorporating various experimental errors. (F) Measured two-point correlation functions of the ground state at half-filling for the regular and (G) disordered lattices, the results for the latter are averaged over $15$ random configurations.
@@
Figure 4: Many-body microwave spectroscopy and temporal decay of quantum spins. (A) Measured dependence of $p$ state occupancies for edge sites (open red and blue circles) and bulk sites (filled gray circles) on microwave detuning $\Delta_{\mathrm{\mu w}}$ at $W=0.4d$. A pronounced peak is observed near zero energy for both edge sites, indicating degeneracy between the two ground states at half-filling. The slight shift of the peak away from $s$ to $p$ resonance is due to van der Waals interactions, which lift the degeneracy between the ground states with $N-1$ and $N$ excitations. (B) Measured magnetization $\sigma_{i}^{z}(t)$ for all 14 sites of the initial Néel state evolved by the realized Hamiltonian in disordered lattices (lower panel) compared to the regular case (upper panel). Rapid loss of magnetization is observed for all sites except the topologically protected edge states of the disordered lattice, whose decays are significantly slower. (C) Empty circles denote results for two edge sites in (B) compared to numerical predictions incorporating various errors in solid lines at $W=0.4{d}$. Filled circles correspond to the two edge states in the regular lattice. Error bars represent SEMs averaged over 15 random configurations. The experimental results for disordered lattices are averaged over $15$ independent realizations.
Structural disorder induced topological phase for many interacting particles.—We next present results for the many-body SPT phase induced by structural disorder at half-filling. Including only nearest-neighbor hoppings and setting $\hat{H}_{\mathrm{vdW}}=0$, the model in Eq. (1) is mapped to a free fermion model via the inverse Jordan-Wigner transformation. Hoppings ($J_{ij}\propto r_{ij}^{-3}$) beyond the nearest neighbors, however, introduce interactions between fermions 26. The van der Waals interactions further give rise to Ising terms, contributing to additional fermion interactions (see SI). The experimentally realized model $\hat{H}$ thus represents a genuine interacting many-body system. Corroborating with the sharp change in the topological invariant (Fig. 1C), the structural disorder induced bosonic average SPT phase also reveals a notable change in the many-body energy spectrum for the open boundary condition case studied. Setting $W=0$, the spectrum features a single ground state, whereas in the disordered case with $W\neq 0$, it exhibits two statistically degenerate ground states (see Fig. 3B and 3C, and SI), which are associated with the occupation of one particle or a $p$ state excitation at either the left or right edge, respectively. Although van der Waals interactions do not break the topology, they nevertheless introduce minor energy shifts at the edges, and cause ground states in the subspaces of $N-1$ or $N+1$ particles to exhibit a small energy difference ($\sim 0.1$ MHz) to those at half-filling (more details in SI).
Many-body ground states with particle numbers near half-filling are prepared experimentally to demonstrate the induced interacting SPT phase. Atoms in an initial product state $\left|\psi_{0}\right\rangle$, e.g., $|1010...10\rangle$, $|0010...10\rangle$, or $|1010...11\rangle$, respectively in the subspaces with $N$, $N-1$, and $N+1$ particles, are arrived in steps: First, all atoms are transferred from the ground $g$ state to the $s$ state via STIRAP, followed by a second adiabatic protocol to the $p$ state from a microwave pulse sweeping across the resonance; focused pinning laser beams are subsequently applied to introduce a spatially dependent AC Stark shift of $\Delta E/h\approx 20$ MHz (Fig. 3A) to the $s$ state; finally, a second microwave pulse is swept across the resonance to produce the product state in which atoms on sites with pinning laser beams remain in the $p$ state while getting de-excited to the $s$ state elsewhere (Fig. 3A).
The product state $\left|\psi_{0}\right\rangle$ evolves into the many-body ground state following the adiabatic ramping down of the pinning laser beams (Fig. 3A), in a process described by the Hamiltonian $\hat{H}_{\mathrm{tot}}=\hat{H}-\Delta E(t)\sum_{i\in S}\hat{n}_{i}$, where $S$ denotes the set of sites with applied pinning lasers. Since this Hamiltonian respects $U(1)$ symmetry, the total particle number is conserved, and the evolving state remains within the subspace of fixed total particle number given by the initial state $\left|\psi_{0}\right\rangle$. At $t=0$, the level shift is significantly larger than the energy scale of hopping terms. As a result, $\left|\psi_{0}\right\rangle$ is a good approximation to the ground state of $\hat{H}_{\mathrm{tot}}$. After a 4 $4\text{ }\mathrm{\mu s}$ pulse of exponential ramping down according to $\Delta E(t)=\Delta Ee^{-t/\tau}$ for $\tau=0.6\text{ }\mathrm{\mu s}$, we arrive at the ground state. Numerically, the above steps achieve a fidelity better than $70\%$ without accounting for errors (see SI).
@@
Figures 3D and 3E present the measured occupancies for all sites of the prepared ground state in both regular and disordered lattices. While their bulk properties are comparable, their behaviors at the edge sites are significantly different. At half-filling, the edge sites in the regular lattice behave similarly to the bulk sites, whereas the disordered configuration reveals pronounced differences for either the left or right edges, as well as the imbalance between the left and right edges. In particular, two statistically degenerate ground states exist in the half-filling subspace, one of which is presented, and is associated with a higher particle occupation at the left edge over the right one. We also observe that in the subspace with $N-1$ particles, the occupancies of edge sites are significantly lower than in the bulk. Conversely, in the subspace with $N+1$ particles, the edge site occupancies are notably higher. These behaviors are consistent with the expected properties of the induced topological ground states.
@@
Ground state degeneracy and quench dynamics.—We now show the observed degeneracy of the ground states in the disordered case. We first prepare a ground state with $N-1$ particles and then employ microwave spectroscopy (as in the single-particle case) with a Rabi frequency of $\Omega_{\mathrm{\mu w}}/(2\pi)\approx 0.25$ MHz and a pulse duration of 2 $2\text{ }\mathrm{\mu s}$ to probe possible excitations by scanning the detuning $\Delta_{\mathrm{\mu w}}$. Figure 4A shows a sharp rise near zero detuning, with a distinct peak appearing at $\Delta_{\mathrm{\mu w}}/(2\pi)\approx-0.1$ MHz for both edge sites, heralding the transition into the half-filling ground state, by adding a particle from absorption of a resonant microwave photon to create a $p$ state atom from an $s$ state at the detuning of $\sim-0.1$ MHz for both edge sites.
Finally, we provide compelling evidence that the phase induced by structural disorder exhibits robust edge states or topological protection, even for highly excited states 52, 53, 54. Specifically, we initialize the system into a product state such as the Néel state $\left|0101\dots 01\right\rangle$, a highly excited state, and subsequently evolve it under the many-body Hamiltonian $\hat{H}$ as in a quench. In Fig. 4B, we observe that in the regular lattice, the magnetization $\sigma_{i}^{z}$ for both bulk and edge sites rapidly decays to zero, implicating a quick loss of initial information or correlation. The small residual value of $\langle\sigma_{i}^{z}\rangle$ is attributed to van der Waals interactions (see SI). In contrast, in the disordered lattice, while the bulk sites exhibit a similar decay to the regular lattice, at the edges $\langle\sigma_{i}^{z}\rangle$ exhibits a slower decay and eventually stabilizes at finite values (Figs. 4B and 4C), in near complete agreement with numerical simulations.
Discussion and outlook.—Our work opens a broad avenue to study the interplay between structural disorder and topology in a fully tunable quantum platform. At the single-particle level, future explorations could address how structural disorder affects higher-dimensional topological phases, such as the Chern insulators 55, 56 and higher-order topological phases 57, 34. At the many-body level with interactions, it is now possible to examine many-body topological phenomena in higher dimensions, including spin liquids in two-dimensional amorphous lattices 58, 59. Furthermore, the Rydberg atom array platform can be employed to explore amorphous Floquet topological phases 60 by incorporating periodic driving. In addition, it also offers new opportunities to study the many-body localization induced by structural disorder 61 and the interplay between localization and topology, such as topologically protected quantum dynamics in disordered systems 52.
Note added. During the completion of this work, we became aware of related work studying average symmetry protected topology in an optical lattice system 62.
@@ Acknowledgements.
We acknowledge significant help and enlightening discussions with Cheng Chen and Songtao Huang, and helpful contributions by Yuanjiang Tang, Chao Liang, Xiangliang Li, Xiaoling Wu, Boyang Wang, and Yifan Wang in the early stages of building up the experimental platform. We acknowledge ChatGPT for language polishing. DMRG calculations were performed using the TeNPy tensor network library 63. We finally acknowledge the support by Center of High Performance Computing, Tsinghua University. This work is supported by the Innovation Program for Quantum Science and Technology (2021ZD0302100). LY is also supported by NSFC (Grants No. 12361131576 and No. 92265205). MKT is supported by NSFC (Grants. No.12234012 and W2431002). YFM, KL, and YX are supported by NSFC (Grants. No.12474265 and 11974201) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301604).
@@
[…]
@@
[…]
@@
Next, the magnetic field is adiabatically increased to approximately 30 Gauss. The optical tweezers are then switched off, and the atoms are excited to the Rydberg $\left|s\right\rangle$ state via two-photon stimulated Raman adiabatic passage (STIRAP) using 420 nm and 1013 nm lasers respectively coupled resonantly to the transitions $g\rightarrow e$ and $e\rightarrow s$. A microwave field at 22.1 GHz is subsequently applied to drive $\left|s\right\rangle$ and $\left|p\right\rangle$ states. Owing to the extremely large polarizability of Rydberg atoms, even a weak microwave field can induce a strong coupling. In our setup, a $-20$ dBm microwave field is sufficient to achieve a Rabi frequency of 2.5 MHz.
@@
At the end of the above sequence, we measure magnetization $\sigma_{i}^{z}$ at each site. Since $[\hat{H},\sigma_{i}^{z}]\neq 0$, it is necessary to pause the interaction during readout. To achieve this, we freeze the system using a 10.38 GHz microwave pulse which excites atoms in $\left|p\right\rangle$ to the $n=53$ hydrogenic manifold, in which atoms experience negligible interactions with those remaining in $\left|s\right\rangle$. The pulse duration is 100 ns, which is sufficiently short to minimize unwanted interactions. Subsequently, a de-excitation pulse is applied to transfer atoms in $\left|s\right\rangle$ back to the ground state using a 1013 nm laser resonantly coupling $\left|s\right\rangle$ to the short-lived intermediate $\left|e\right\rangle$ state from which atoms decay back to the ground state.
@@
We compare the experimental results with simulations considering different imperfections in the main text. Here we discuss the main source of errors in our system, including state preparation, detection errors (SPAM), and those due to thermal motion of atoms.
@@
while the Rydberg $\left|p\right\rangle$ state remains unaffected as it is not coupled to the short lived $\left|e\right\rangle$ state by this addressing laser. To mitigate this detrimental effect of the addressing laser, we choose first to transfer all atoms into the Rydberg $\left|p\right\rangle$ state; during this process, the presence of the dipolar and van der Waals interactions gives rise to an error $\eta_{\mathrm{MW}}\approx 0.02$. The addressing lasers are then applied to sub-chain $\alpha$, inducing an AC Stark shift of approximately 20 MHz for all $\left|s\right\rangle$ state atoms in sub-chain $\alpha$. Finally we sweep the MW frequency across the single-atom resonance to transfer atoms in sub-chain $\beta$ to $\left|s\right\rangle$ states, with an associated error $\eta_{\beta}\approx 0.05$, while atoms in sub-chain $\alpha$ remain in $\left|p\right\rangle$ state with an error $\eta_{\alpha}\approx 0.04$.
@@
$[H^{\textrm{S}}]_{ij}=J_{ij}(1-\delta_{ij})-V_{i}^{\text{vdW}}\delta_{ij},$ (S2)
@@
Without van der Waals interactions ($V_{i}^{\text{vdW}}=0$), the single-particle Hamiltonian in Eq. (S2) respects the sublattice symmetry, i.e., $\Pi H^{\textrm{S}}\Pi^{-1}=-H^{\textrm{S}}$, with $\Pi=\text{diag}\{(-1)^{j}\}_{j=1}^{2N}$ the sublattice symmetry operator. Although the $\mathbb{Z}$ classification applies for the topology protected by sublattice symmetry 3, our model exhibits only two phases: a topologically trivial phase without edge modes and a nontrivial phase with one edge state at each edge. As such, we can use the polarization, which is equivalent to the Berry phase in a regular lattice, defined as 47
@@
to characterize the topology of the Hamiltonian. Here $U=(|u_{1}\rangle,|u_{2}\rangle,\dots,|u_{N}\rangle)$, with $|u_{j}\rangle$ ($j=1,2,\dots N$) denoting the lowest $N$ eigenstates of the single-particle Hamiltonian $H^{\mathrm{S}}$ under periodic boundary conditions, and $D=\mathrm{diag}\{e^{2\pi ix_{j}/(N{d})}\}_{j=1}^{2N}$, with ${x}_{j}$ the spatial coordinate of the $\lfloor(j+1)/2\rfloor$th unit cell. It is well known that with the sublattice symmetry, $P_{\mathrm{S}}$ can only take the values of either 00 or $0.5$ 36. This restriction allows $P_{\mathrm{S}}$ to function as a topological invariant: For a topologically trivial phase, $P_{\mathrm{S}}=0$, while for a nontrivial phase, $P_{\mathrm{S}}=0.5$.
@@
Previous studies have shown that the inversion symmetry can protect the topology in 1D 64, 65. We now present a proof showing that, under this symmetry, the polarization $P_{\mathrm{S}}$ can only take the values of either 00 or $0.5$. Specifically, consider the Hamiltonian $H^{\textrm{S}}$ that respects the inversion symmetry, i.e., $[H^{\textrm{S}},U_{R}]=0$, where $U_{R}$ is the inversion operator. When acting on $|i\rangle$, the operator yields $U_{R}|i\rangle=|\mathcal{R}(i)\rangle$, where $\mathcal{R}(i)$ denotes the site index of the inversion partner of site $i$. This symmetry ensures that $|u_{i}\rangle$ in $U$ is also an eigenstate of $U_{R}$ corresponding to an eigenvalue $\nu_{i}$ that takes a value of $1$ or $-1$. With this in mind, we can derive that
@@
Substituting into Eq. (S4) and subsequently into Eq. (S3), then we obtain
@@
Let $\det(U^{\dagger}DU)=re^{i2\pi\theta}$ ($r>0$). Then we obtain $P_{\mathrm{S}}=(\theta-x_{c}/{d})\textrm{ mod }1=(-\theta+x_{c}/{d})\textrm{ mod }1$, yielding $\theta=x_{c}/{d}+n/2$ for integer $n$. Therefore, $P_{\mathrm{S}}=n/2\textrm{ mod }1$ is quantized to 00 or $0.5$ by the inversion symmetry.

### v3

@@ Abstract
Figure 3: Observation of structural disorder-induced many-body average SPT phase. a, Schematics of the experimental sequence for many-body ground state preparation and detection. Initially, all atoms are excited into the Rydberg $p$ state (top left panel). Focused pinning laser beams as in Fig.1 are applied to atoms in the upper sub-chain, and a microwave field is swept across resonance to produce a Néel state (right panel). Adiabatic evolution subsequently prepares the ground state of the Hamiltonian $\hat{H}$ in a fixed particle number subspace. b, Energy spectra at different particle numbers for a regular lattice and c, disordered lattice (W=0.4dW=0.4{d}), with a zoomed-in view near the ground states. The four-fold degeneracy of the ground state is slightly lifted due to the van der Waals interactions. The average inversion symmetry only protects the two-fold ground state degeneracy in the half-filling subspace (see the SI). d, Measured site occupancy of the ground state for the regular lattice compared to (e) the disordered lattice near half-fillings of N1N-1 (left), NN (right), and $N+1$ (below) particles. The measured results are postselected for specific particle fillings. Diagonal crosses represent numerical results incorporating various experimental errors. Measured two-point correlation functions of the ground state at half-filling for (f) the regular and (g) disordered lattices.
@@
Figure 4: Many-body microwave spectroscopy and temporal decay of quantum spins. a, Measured dependence of $p$ state occupancies for edge sites (open red and blue circles) and bulk sites (filled gray circles) on microwave detuning $\Delta_{\mathrm{\mu w}}$ at $W=0.4d$. A pronounced peak is observed near zero energy for both edge sites, indicating degeneracy between the two ground states at half-filling. The slight shift of the peak away from $s$ to $p$ resonance is due to van der Waals interactions, which lift the degeneracy between the ground states with N1N-1 and NN excitations. b, Measured magnetization $\sigma_{i}^{z}(t)$ for all 14 sites of the initial Néel state evolved by the realized Hamiltonian in disordered lattices (lower panel) compared to the regular case (upper panel). Rapid loss of magnetization is observed for all sites except the topologically protected edge states of the disordered lattice, whose decays are significantly slower. c, Empty circles denote results for two edge sites in b compared to numerical predictions incorporating various errors in solid lines at W=0.4dW=0.4{d}. Filled circles correspond to the two edge states in the regular lattice. Error bars represent SEMs averaged over 15 random configurations.
We next present results for the many-body interacting average SPT phase induced by structural disorder at half-filling. Many-body ground states with particle numbers near half-filling are prepared experimentally to demonstrate the induced SPT phase. Initially, atoms are prepared in a product state $\left|\psi_{0}\right\rangle$, e.g., $|1010...10\rangle$, $|0010...10\rangle$, or $|1010...11\rangle$, respectively in the subspaces with $N$, N1N-1, and $N+1$ particles, as illustrated Fig. 3a and detailed in the SI. The product state $\left|\psi_{0}\right\rangle$ then evolves into the many-body ground state following the adiabatic ramping down of the pinning laser beams (see Fig. 3a and SI).
@@
Figures 3d and 3e present the measured occupancies for all sites of the prepared ground state in both regular and disordered lattices. While their bulk properties are comparable, their behaviors at the edge sites are significantly different. At half-filling, the edge sites in the regular lattice behave similarly to the bulk sites consistent with the fact that the spectrum features a single ground state (see Fig. 3b), whereas the disordered configuration reveals pronounced differences for either the left or right edges, as well as the imbalance between the left and right edges. In particular, two statistically degenerate ground states exist in the half-filling subspace (see Fig. 3c and SI), one of which is presented, and is associated with a higher particle occupation at the left edge over the right one. We also observe that in the subspace with N1N-1 particles, the occupancies of edge sites are significantly lower than in the bulk. Conversely, in the subspace with $N+1$ particles, the edge site occupancies are notably higher. These behaviors are consistent with the expected properties of the induced topological ground states.
@@
We now show the observed statistical degeneracy of the ground states in the half-filling subspace in the disordered case. We first prepare a ground state with N1N-1 particles and then employ microwave spectroscopy (as in the single-particle case) with a Rabi frequency of $\Omega_{\mathrm{\mu w}}/(2\pi)\approx 0.25$ MHz and a pulse duration of $2\text{ }\mathrm{\mu s}$ to probe possible excitations by scanning the detuning $\Delta_{\mathrm{\mu w}}$. Figure 4a shows a sharp rise near zero detuning, with a distinct peak appearing at $\Delta_{\mathrm{\mu w}}/(2\pi)\approx-0.1$ MHz for both edge sites, heralding the transition into the half-filling ground state, by adding a particle from absorption of a resonant microwave photon to create a $p$ state atom from an $s$ state at the detuning of $\sim-0.1$ MHz for both edge sites. Here, the small detuning arises from van der Waals interactions (more details in SI).
Finally, we provide compelling evidence that the disorder-induced phase exhibits robust edge states or topological protection, even for highly excited states 65, 66, 67. Specifically, we initialize the system into a product state such as the Néel state $\left|0101\dots 01\right\rangle$, a highly excited state, and subsequently evolve it under the many-body Hamiltonian $\hat{H}$ as in a quench. In Fig. 4b, we observe that in the regular lattice, the magnetization $\sigma_{i}^{z}$ for both bulk and edge sites rapidly decays to zero, implicating a quick loss of initial information. The small residual value of $\langle\sigma_{i}^{z}\rangle$ is attributed to van der Waals interactions (see SI). In contrast, in the disordered lattice, while the bulk sites exhibit a similar decay to the regular lattice, at the edges $\langle\sigma_{i}^{z}\rangle$ exhibits a slower decay and eventually stabilizes at finite values (Figs. 4b and 4c), in near complete agreement with numerical simulations.
Our work opens a broad avenue to study the interplay between structural disorder and topology in a programable quantum platform. At the single-particle level, future explorations could address how structural disorder affects higher-dimensional topological phases, such as the Chern insulators 68, 69 and higher-order topological phases 70, 15. At the many-body level with interactions, it is now possible to examine many-body topological phenomena in higher dimensions, including spin liquids in two-dimensional amorphous lattices 71, 72. Furthermore, the Rydberg atom array platform can be employed to explore amorphous Floquet topological phases 73 by incorporating periodic driving. In addition, it also offers new opportunities to study the many-body localization induced by structural disorder 74 and the interplay between localization and topology, such as topologically protected quantum dynamics in disordered systems 65.
Note added. During the completion of this work, we became aware of related experiment on average symmetry protected topology in an optical lattice system 75.
@@ Acknowledgements.
We acknowledge significant help and enlightening discussions with Cheng Chen and Songtao Huang, and helpful contributions by Yuanjiang Tang, Chao Liang, Xiangliang Li, Xiaoling Wu, Boyang Wang, and Yifan Wang in the early stages of building up the experimental platform. We acknowledge ChatGPT for language polishing. DMRG calculations were performed using the TeNPy tensor network library 76. We finally acknowledge the support by Center of High Performance Computing, Tsinghua University. This work is supported by the Innovation Program for Quantum Science and Technology (2021ZD0302100). LY is also supported by NSFC (Grants No. 12361131576 and No. 92265205). MKT is supported by NSFC (Grants. No.12234012 and W2431002). YFM, KL, and YX are supported by NSFC (Grants. No.12474265 and 11974201) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301604).
@@
The realization of the model in Eq. (1) in the main text is based on a 87Rb Rydberg atom array. We encode Rydberg state $\left|s\right\rangle$ = $|55S_{1/2}\rangle$ as $|\uparrow\rangle$ and $\left|p\right\rangle$ = $|55P_{1/2}\rangle$ as $|\downarrow\rangle$. The two pseudo-spin states are coupled by a resonant microwave at 22.1 GHz. A 30 Gauss magnetic field, aligned parallel to the plane formed by the two sub-chains, is applied to split different Zeeman states.
@@
The initial product states are prepared with the help of an addressing laser, split from the 1013 nm laser beam and detuned from the transition between the $|\uparrow\rangle$ and the intermediate excited state $\left|e\right\rangle$ = $|6P_{3/2}\rangle$. The addressing laser is locked to the stable 1013 nm laser employed in two-photon excitation to Rydberg states via their beat frequency, allowing for its frequency to be tuned and is set 350 MHz above the resonance.
@@
Next, the magnetic field is adiabatically increased to approximately 30 Gauss. The optical tweezers are then switched off, and the atoms are excited to the Rydberg $\left|s\right\rangle$ state via two-photon stimulated Raman adiabatic passage (STIRAP) using 420 nm and 1013 nm lasers respectively coupled resonantly to the transitions $g\rightarrow e$ and $e\rightarrow s$. A microwave field at 22.1 GHz is subsequently applied to drive $\left|s\right\rangle$ and $\left|p\right\rangle$ states. Owing to the extremely large polarizability of Rydberg atoms, even a weak microwave field can induce a strong coupling. In our setup, a 20-20 dBm microwave field is already sufficient to achieve a Rabi frequency of 2.5 MHz.
@@
At the end of the above sequence, we measure magnetization $\sigma_{i}^{z}$ at each site. Since $\hat{H},\sigma_{i}^{z}]\neq 0$, it is necessary to pause the interaction during readout. To achieve this, we freeze the system using a 10.38 GHz microwave pulse which excites atoms in $\left|p\right\rangle$ to the $n=53$ hydrogenic manifold, in which atoms experience negligible interactions with those remaining in $\left|s\right\rangle$. The pulse duration is 100 ns, which is sufficiently short to minimize unwanted interactions. Subsequently, a de-excitation pulse is applied to transfer atoms in $\left|s\right\rangle$ back to the ground state using a 1013 nm laser resonantly coupling $\left|s\right\rangle$ to the short-lived intermediate $\left|e\right\rangle$ state from which atoms decay back to the ground state.
I.5 S1.5. Adiabatic preparation of the many-body ground state
We experimentally prepare the many-body ground states with particle numbers near half-filling. Atoms in an initial product state $\left|\psi_{0}\right\rangle$, e.g., $|1010...10\rangle$, $|0010...10\rangle$, or $|1010...11\rangle$, respectively in the subspaces with $N$, N1N-1, and $N+1$ particles, are arrived in steps: First, all atoms are transferred from the ground $g$ state to the $s$ state via STIRAP, followed by a second adiabatic protocol to the $p$ state from a microwave pulse sweeping across the resonance; focused pinning laser beams are subsequently applied to introduce a spatially dependent AC Stark shift of $\Delta E/h\approx 20$ MHz (Fig. 3a in the main text) to the $s$ state; finally, a second microwave pulse is swept across the resonance to produce the product state in which atoms on sites with pinning laser beams remain in the $p$ state while getting de-excited to the $s$ state elsewhere (Fig. 3a in the main text).
The product state $\left|\psi_{0}\right\rangle$ evolves into the many-body ground state following the adiabatic ramping down of the pinning laser beams (Fig. 3a in the main text), in a process described by the Hamiltonian
$\hat{H}_{\mathrm{tot}}=\hat{H}-\Delta E(t)\sum_{i\in S}\hat{n}_{i},$ (S1)
where $S$ denotes the set of sites with applied pinning lasers. Since this Hamiltonian respects $U(1)$ symmetry, the total particle number is conserved, and the evolving state remains within the subspace of fixed total particle number given by the initial state $\left|\psi_{0}\right\rangle$. At $t=0$, the level shift is significantly larger than the energy scale of hopping terms. As a result, $\left|\psi_{0}\right\rangle$ is a good approximation to the ground state of $\hat{H}_{\mathrm{tot}}$. After a $4\text{ }\mathrm{\mu s}$ pulse of exponential ramping down according to $\Delta E(t)=\Delta Ee^{-t/\tau}$ for $\tau=0.6\text{ }\mathrm{\mu s}$, we arrive at the ground state experimentally.
The effectiveness of the adiabatic evolution depends on the low energy spectral gap. Figure S1 displays the numerically calculated energy gap during the time evolution. We see that for the regular configuration, the gap quickly decreases, reaching a minimum value of $\Delta_{\mathrm{g}}/h=1.06$ MHz. For disordered configurations, the minimum energy gaps averaged over random configurations are $0.75$ MHz, $0.89$ MHz, and $0.72$ MHz for $N$, N1N-1, and $N+1$ particles, respectively, which are slightly smaller than the minimum energy gap in the regular case. The energy gaps are on the order of the energy scale of the hopping term in the Hamiltonian $\hat{H}$. At half-filling, it is defined as the gap between the third lowest energy state with respect to the ground state, as the ground state is nearly two-fold degenerate. During the time evolution, although the gap declines, the ramping speed for the pinning laser beams decreases correspondingly, thereby ensuring a successful adiabatic evolution.
Figure S1 also plots the fidelity between the numerically calculated evolving state and the corresponding true ground state during the time evolution without accounting for errors, where the evolving state is obtained by simulating the time evolution of the Hamiltonian using the Krylov method. We see that at the end of the adiabatic evolution, the fidelity reaches $0.99$ for the regular configuration. For disordered cases, their average fidelity values are $0.89$, $0.75$, and $0.72$ for the fillings of $N$, N1N-1, and $N+1$ particles, respectively. To ensure that imperfections of the experiments are well understood, we also perform simulations by considering all SPAM errors. The results of numerical simulations are found to be in good agreement with the experimental data, as demonstrated in Fig. 3 in the main text.
@@
We compare the experimental results with numerical simulations considering different imperfections in the main text. Here we discuss the main source of errors in our system, including state preparation, detection errors (SPAM), and those due to thermal motion of atoms.
@@
while the Rydberg $\left|p\right\rangle$ state remains unaffected as it is not coupled to the short lived $\left|e\right\rangle$ state by this addressing laser. To mitigate this detrimental effect of the addressing laser, we choose first to transfer all atoms into the Rydberg $\left|p\right\rangle$ state; during this process, the presence of the dipolar and van der Waals interactions gives rise to an error $\eta_{\mathrm{MW}}\approx 0.02$. The addressing lasers are then applied to sub-chain $\alpha$, inducing an AC Stark shift of approximately 20 MHz for all $\left|s\right\rangle$ state atoms in sub-chain $\alpha$. Finally we sweep the MW frequency back across the single-atom resonance to transfer atoms in sub-chain $\beta$ to $\left|s\right\rangle$ states, with an associated error $\eta_{\beta}\approx 0.05$, while atoms in sub-chain $\alpha$ remain in $\left|p\right\rangle$ state with an error $\eta_{\alpha}\approx 0.04$.
@@
$H^{\textrm{S}}]_{ij}=J_{ij}(1-\delta_{ij})-V_{i}^{\text{vdW}}\delta_{ij},$ (S3)
@@
Without van der Waals interactions ($V_{i}^{\text{vdW}}=0$), the single-particle Hamiltonian in Eq. (S3) respects the sublattice symmetry, i.e., $\Pi H^{\textrm{S}}\Pi^{-1}=-H^{\textrm{S}}$, with $\Pi=\text{diag}\{(-1)^{j}\}_{j=1}^{2N}$ the sublattice symmetry operator. Although the $\mathbb{Z}$ classification applies for the topology protected by sublattice symmetry 58, our model exhibits only two phases: a topologically trivial phase without edge modes and a nontrivial phase with one edge state at each edge. As such, we can use the polarization, which is equivalent to the Berry phase in a regular lattice, defined as 59
@@
to characterize the topology of the Hamiltonian. Here $U=(|u_{1}\rangle,|u_{2}\rangle,\dots,|u_{N}\rangle)$, with $|u_{j}\rangle$ ($j=1,2,\dots N$) denoting the lowest $N$ eigenstates of the single-particle Hamiltonian $H^{\mathrm{S}}$ under periodic boundary conditions, and $D=\mathrm{diag}\{e^{2\pi ix_{j}/(N{d})}\}_{j=1}^{2N}$, with ${x}_{j}$ the spatial coordinate of the $\lfloor(j+1)/2\rfloorth$ unit cell. It is well known that with the sublattice symmetry, $P_{\mathrm{S}}$ can only take the values of either 0 or $0.5$ 39, which allows $P_{\mathrm{S}}$ to function as a topological invariant: For a topologically trivial phase, $P_{\mathrm{S}}=0$, while for a nontrivial phase, $P_{\mathrm{S}}=0.5$.
@@
Previous studies have shown that the inversion symmetry can protect the topology in 1D 77, 78. We now present a proof showing that, under this symmetry, the polarization $P_{\mathrm{S}}$ can only take the values of either 0 or $0.5$. Specifically, consider the Hamiltonian $H^{\textrm{S}}$ that respects the inversion symmetry, i.e., $H^{\textrm{S}},U_{R}]=0$, where $U_{R}$ is the inversion operator. When acting on $|i\rangle$, the operator yields $U_{R}|i\rangle=|\mathcal{R}(i)\rangle$, where $\mathcal{R}(i)$ denotes the site index of the inversion partner of site $i$. This symmetry ensures that $|u_{i}\rangle$ in $U$ is also an eigenstate of $U_{R}$ corresponding to an eigenvalue $\nu_{i}$ that takes a value of $1$ or 1-1. With this result, we can derive
@@
Substituting into Eq. (S5) and subsequently into Eq. (S4), we then obtain
@@
Let $\det(U^{\dagger}DU)=re^{i2\pi\theta}$ ($r>0$). Then we obtain $P_{\mathrm{S}}=(\theta-x_{c}/{d})\textrm{ mod }1=(-\theta+x_{c}/{d})\textrm{ mod }1$, yielding $\theta=x_{c}/{d}+n/2$ for integer $n$. Therefore, $P_{\mathrm{S}}=n/2\textrm{ mod }1$ is quantized to 0 or $0.5$ by the inversion symmetry.